英語 (English)

主題別英語 (Thematic English) (工 ((知 B)1 年))

ニムチャックアーレン · 非常勤講師 1 単位後期木 1 ・ 2

(平成 19 年度以前の授業科目: 『主題別英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】 Students will work on practicing discussion. The goal is for students

to be active, not passive in class discussions. 【授業の概要】Students will discuss a different topic each week, practicing the

English necessary for each week’s lesson. 【キーワード】 discussion; communication 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】 Students should come to class and try to use only English, without

fear of making mistakes. Mistakes will happen, and students should not worry when they make them. Enjoy the class, have fun, and use English.

【授業の計画】 1. Introductions

2. Bookcrossing

3. Names

4. Gender roles

5. Movies

6. Food for thought

7. Winter

8. Mid-term Test

9. Sleep and Dreams

10. Stress

11. Holidays

12. Opinions

13. Weather

14. Gestures

15. Canada

16. Final test 【教科書】無【参考書等】 Practical English Textbook - Thematic (\begin{matrix}Y \\ = \end{matrix}500) - available from teacher,

2nd week 【成績評価の方法】 Students’ grades will be as follows:Attendance = 50% of students’ grades.Mid- term test = 25% of students’ grades. Final test = 25% of students’ grades.Final grades will consist of an attendance score and test scores. (ie. 50% + 25% + 25% = 100%)

【再試験の有無】無

【受講者へのメッセージ】 Have fun, enjoy and speak English!

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221192 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】

⇒ ニムチャック . (オフィスアワー: canadaclubjp@bloom.ocn.ne.jp)

English

Thematic English (工 ((知 B)1 年))

ARLEN NIMCHUK · Part-time Lecturer 1 unit 後期木 1 ・ 2

(平成 19 年度以前の授業科目: 『主題別英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ Students will work on practicing discussion. The goal is for students to

be active, not passive in class discussions. Outline⟩ Students will discuss a different topic each week, practicing the English

necessary for each week’s lesson. �������⟩ discussion; communication Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ Students should come to class and try to use only English, without fear of

making mistakes. Mistakes will happen, and students should not worry when they make them. Enjoy the class, have fun, and use English.

Schedule⟩

1. Introductions

2. Bookcrossing

3. Names

4. Gender roles

5. Movies

6. Food for thought

7. Winter

8. Mid-term Test

9. Sleep and Dreams

10. Stress

11. Holidays

12. Opinions

13. Weather

14. Gestures

15. Canada

16. Final test Textbook⟩ 無 Reference⟩ Practical English Textbook - Thematic (\begin{matrix}Y \\ = \end{matrix}500) - available from teacher,

2nd week Evaluation Criteria⟩ Students’ grades will be as follows:Attendance = 50% of students’ grades.Mid- term test = 25% of students’ grades. Final test = 25% of students’ grades.Final grades will consist of an attendance score and test scores. (ie. 50% + 25% + 25% = 100%)

Re-evaluation⟩ 無

Message⟩ Have fun, enjoy and speak English!

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221192 Contact (Office-Hour, Room, E-mail)⟩

⇒ NIMCHUK . (Office Hour: canadaclubjp@bloom.ocn.ne.jp)